# Learning Adaptive Pseudo-Label Selection for Semi-Supervised 3D Object Detection

### v2

@@ Abstract
Semi-supervised 3D object detection (SS3DOD) aims to reduce costly 3D annotations utilizing unlabeled data. Recent studies adopt pseudo-label-based teacher-student frameworks and demonstrate impressive performance. The main challenge of these frameworks is in selecting high-quality pseudo-labels from the teacher’s predictions. Most previous methods, however, select pseudo-labels by comparing confidence scores over thresholds manually set. The latest works tackle the challenge either by dynamic thresholding or refining the quality of pseudo-labels. Such methods still overlook contextual information e.g. object distances, classes, and learning states, and inadequately assess the pseudo-label quality using partial information available from the networks. In this work, we propose a novel SS3DOD framework featuring a learnable pseudo-labeling module designed to automatically and adaptively select high-quality pseudo-labels. Our approach introduces two networks at the teacher output level. These networks reliably assess the quality of pseudo-labels by the score fusion and determine context-adaptive thresholds, which are supervised by the alignment of pseudo-labels over GT bounding boxes. Additionally, we introduce a soft supervision strategy that can learn robustly under pseudo-label noises. This helps the student network prioritize cleaner labels over noisy ones in semi-supervised learning. Extensive experiments on the KITTI and Waymo datasets demonstrate the effectiveness of our method. The proposed method selects high-precision pseudo-labels while maintaining a wider coverage of contexts and a higher recall rate, significantly improving relevant SS3DOD methods.
@@ 1 Introduction
The pseudo-label-based framework has been most widely adopted for Semi-Supervised 3D Object Detection (SS3DOD), effectively leveraging unlabeled data. Variants of SS3DOD methods [40, 52, 9, 56, 20, 12, 6, 16, 44, 41] based on this framework have been developed, achieving significant performance gains. Within this framework, the pseudo-labeling plays a critical role in detection performance. Previous studies typically select pseudo-labels by thresholding scores predicted by the teacher network (e.g. classification confidence and objectness). For instance, [40, 9, 20, 41, 16, 44] manually set the thresholds to filter pseudo-labels. ATF-3D [56] introduced a threshold searching mechanism by object distances and classes. The state-of-the-art method, HSSDA [12], automatically set dual thresholds. For three object classes, nine clustering steps on three scores determine 18 thresholds for hierarchical supervision. Using these thresholds directly on the teacher’s scores causes suboptimal pseudo-label selection for two reasons. First, predicting the quality of pseudo-labels is challenging. In SS3DOD, the base detector typically yields multiple scores. Each score exhibits a different correlation with GT pseudo-label quality (see Fig. 4). This complicates setting a consistent threshold across multiple scores and assessing the reliability of pseudo-labels. As a result, previous methods relied only on partial information rather than utilizing all available indicators for assessing pseudo-label quality (see Fig. 2(d)). Second, the optimal threshold must account for the context of instances (e.g. object classes, distances, and learning states). Fig. 2 illustrates that the score distributions vary across classes, distances, and learning states. The optimal threshold is consequently context-dependent. Rather than a fixed threshold, an effective threshold needs to balance the quality and coverage of pseudo-labels across contexts. Moreover, addressing the dynamic learning states during training requires recalculating the threshold for SSL. Finding an optimal threshold that accounts for such contexts is complex, approaches such as [40, 56, 9, 12] only partially consider these contextual factors (see Fig. 2(d))
@@
To address the aforementioned limitations of prior arts, we propose a novel learning-based pseudo-label selection method, named Pseudo-label Selection Module (PSM). The PSM leverages limited Ground Truth (GT) information to assess the quality of pseudo-labels and to determine a context-appropriate threshold. The PSM consists of the Pseudo-Label Quality Estimator (PQE) and Context-aware Threshold Estimator (CTE). The PQE encodes the teacher’s various output scores to a single fusion score indicating reliable pseudo-label quality, while the CTE encodes the context to generate adaptive threshold values. During SSL, PSM is trained to dynamically select pseudo-labels considering the context, achieving a wide coverage of pseudo-labels while maintaining a high quality. Additionally, we introduce the Soft Supervision strategy to train robustly against pseudo-label noises. Our method combines the soft GT sampling augmentation and loss re-weighting to counteract pseudo-label noises given the coverage of labels. To the best of our knowledge, this is the first method to model pseudo-labeling using a neural network. Our contributions are summarized as follows:
@@ 2 Related Work
3D object detection involves predicting oriented 3D bounding boxes and types of objects from from either monocular RGB images [31, 32, 4, 33, 48] or LiDAR point clouds. 3D object detection from LiDAR point clouds is generally categorized into point-based detectors [26, 30, 50, 51, 23, 17] and grid-based detectors [49, 28, 2, 53, 14, 7, 43] by data representations. Point-based detectors encode spatial information directly from the raw point cloud. In PointRCNN [26], the point-based backbone [22] hierarchically samples the input point cloud and extracts features. The extracted point-level features are used in a two-stage process to generate and refine 3D proposals. In contrast, grid-based detectors convert the sparse point cloud into a grid representation (e.g. voxels and pillars) and then efficiently encode the scene using conventional CNNs. VoxelNet [59] extracts features from points within voxels using PointNet [21], and then generates 3D detection proposals using 3D CNNs. SECOND [49] improves speed and efficiency by applying sparse 3D convolutions instead of dense 3D convolutions. Additionally, [7, 43] stacks the point cloud into vertical columns (pillars) and applies a simplified PointNet to extract features from each pillar. Point-voxel-based detectors combine the strengths of both grid-based and point-based detectors. Studies such as [27, 29, 13, 11, 25] utilize voxel- and point-based operations for 3D proposal generation or refinement. In this work, we conducted experiments using the grid-based detector Voxel-RCNN [2] and the point-voxel-based detector PV-RCNN [27].

### v3

@@ Abstract
Semi-supervised 3D object detection (SS3DOD) aims to reduce costly 3D annotations utilizing unlabeled data. Recent studies adopt pseudo-label-based teacher-student frameworks and demonstrate impressive performance. The main challenge of these frameworks is in selecting high-quality pseudo-labels from the teacher’s predictions. Most previous methods, however, select pseudo-labels by comparing confidence scores over thresholds manually set. The latest works tackle the challenge either by dynamic thresholding or refining the quality of pseudo-labels. Such methods still overlook contextual information e.g., object distances, classes, and learning states, and inadequately assess the pseudo-label quality using partial information available from the networks. In this work, we propose a novel SS3DOD framework featuring a learnable pseudo-labeling module designed to automatically and adaptively select high-quality pseudo-labels. Our approach introduces two networks at the teacher output level. These networks reliably assess the quality of pseudo-labels by the score fusion and determine context-adaptive thresholds, which are supervised by the alignment of pseudo-labels over GT bounding boxes. Additionally, we introduce a soft supervision strategy that can learn robustly under pseudo-label noise. This helps the student network prioritize cleaner labels over noisy ones in semi-supervised learning. Extensive experiments on the KITTI and Waymo datasets demonstrate the effectiveness of our method. The proposed method selects high-precision pseudo-labels while maintaining a wider coverage of contexts and a higher recall rate, significantly improving relevant SS3DOD methods.
@@ 1 Introduction
The pseudo-label-based framework has been most widely adopted for Semi-Supervised 3D Object Detection (SS3DOD), effectively leveraging unlabeled data. Variants of SS3DOD methods [40, 52, 9, 56, 20, 12, 6, 16, 44, 41] based on this framework have been developed, achieving significant performance gains. Within this framework, the pseudo-labeling plays a critical role in detection performance. Previous studies typically select pseudo-labels by thresholding scores predicted by the teacher network (e.g., classification confidence and objectness). For instance, [40, 9, 20, 41, 16, 44] manually set the thresholds to filter pseudo-labels. ATF-3D [56] introduced a threshold searching mechanism by object distances and classes. The state-of-the-art method, HSSDA [12], automatically set dual thresholds. For three object classes, nine clustering steps on three scores determine 18 thresholds for hierarchical supervision. Using these thresholds directly on the teacher’s scores causes suboptimal pseudo-label selection for two reasons. First, predicting the quality of pseudo-labels is challenging. In SS3DOD, the base detector typically yields multiple scores. Each score exhibits a different correlation with GT pseudo-label quality (see Fig. 4). This complicates setting a consistent threshold across multiple scores and assessing the reliability of pseudo-labels. As a result, previous methods relied only on partial information rather than utilizing all available indicators for assessing pseudo-label quality (see Fig. 2(d)). Second, the optimal threshold must account for the context of instances (e.g., object classes, distances, and learning states). Fig. 2 illustrates that the score distributions vary across classes, distances, and learning states. The optimal threshold is consequently context-dependent. Rather than a fixed threshold, an effective threshold needs to balance the quality and coverage of pseudo-labels across contexts. Moreover, addressing the dynamic learning states during training requires recalculating the threshold for SSL. Finding an optimal threshold that accounts for such contexts is complex, approaches such as [40, 56, 9, 12] only partially consider these contextual factors (see Fig. 2(d))
@@
To address the aforementioned limitations of prior work, we propose a novel learning-based pseudo-label selection method, named Pseudo-label Selection Module (PSM). The PSM leverages limited Ground Truth (GT) information to assess the quality of pseudo-labels and to determine a context-appropriate threshold. The PSM consists of the Pseudo-Label Quality Estimator (PQE) and Context-aware Threshold Estimator (CTE). The PQE encodes the teacher’s various output scores to a single fusion score indicating reliable pseudo-label quality, while the CTE encodes the context to generate adaptive threshold values. During SSL, PSM is trained to dynamically select pseudo-labels considering the context, achieving a wide coverage of pseudo-labels while maintaining a high quality. Additionally, we introduce the Soft Supervision strategy to train robustly against pseudo-label noises. Our method combines the soft GT sampling augmentation and loss re-weighting to counteract pseudo-label noises given the coverage of labels. To the best of our knowledge, this is the first method to model pseudo-labeling using a neural network. Our contributions are summarized as follows:
@@ 2 Related Work
3D object detection involves predicting oriented 3D bounding boxes and types of objects from either monocular RGB images [31, 32, 4, 33, 48] or LiDAR point clouds. 3D object detection from LiDAR point clouds is generally categorized into point-based detectors [26, 30, 50, 51, 23, 17] and grid-based detectors [49, 28, 2, 53, 14, 7, 43] by data representations. Point-based detectors encode spatial information directly from the raw point cloud. In PointRCNN [26], the point-based backbone [22] hierarchically samples the input point cloud and extracts features. The extracted point-level features are used in a two-stage process to generate and refine 3D proposals. In contrast, grid-based detectors convert the sparse point cloud into a grid representation (e.g., voxels and pillars) and then efficiently encode the scene using conventional CNNs. VoxelNet [59] extracts features from points within voxels using PointNet [21], and then generates 3D detection proposals using 3D CNNs. SECOND [49] improves speed and efficiency by applying sparse 3D convolutions instead of dense 3D convolutions. Additionally, [7, 43] stacks the point cloud into vertical columns (pillars) and applies a simplified PointNet to extract features from each pillar. Point-voxel-based detectors combine the strengths of both grid-based and point-based detectors. Studies such as [27, 29, 13, 11, 25] utilize voxel- and point-based operations for 3D proposal generation or refinement. In this work, we conducted experiments using the grid-based detector Voxel-RCNN [2] and the point-voxel-based detector PV-RCNN [27].